### 07 Apr 2026 v1

Abstract
[…]
keywords:
cosmology, large scale structure of the Universe, computer modeling and simulation, astronomical catalogs
[…]
98.80.-k, 98.65.-r, 95.75.Pq, 95.80.+p
1 Introduction
[…]
A key quantity in this context is the subhalo merger timescale, which characterizes the time required for an infalling subhalo to sink and coalesce with its host. Several numerical studies have derived fitting functions for this timescale using both idealized and cosmological simulations [9, 10, 11, 12, 13]. These works consistently show that the infall redshift is an important factor, as it sets the dynamical time of the host halo at accretion: subhalos accreted at higher redshifts (earlier times) generally experience shorter dynamical times and thus merge more rapidly. Other critical factors established by these works include the orbital parameters (e.g., circularity) and the subhalo-to-host mass ratio at infall: subhalos on more radial orbits or with higher mass ratios generally merge more rapidly.
Merger timescale prescriptions play a central role in semi-analytic models [14, 15, 16, 17], where they are used to translate the merger history of subhalos into predictions for galaxy–galaxy mergers. This connection enables SAMs to study a range of merger-driven physical processes, such as morphological transformations [18, 19, 20], starbursts [21, 22, 23, 24], and AGN activity [25, 26].
The subhalo merger rate is another key quantity, describing the frequency of merger events in a population [27, 28, 29, 30, 31, 32, 33]. While it is fundamentally defined as the number of mergers occurring per unit time per subhalo, the precise definition varies across studies. For instance, some works measure the rate per descendant halo [27], while others report it per progenitor subhalo [28]. Despite these methodological differences, estimates from different models are found to be broadly consistent, typically agreeing within a factor of two [34].
A key complication in these studies is that there is no universal definition of what a subhalo “merger” is. The lack of a unique definition complicates the interpretation of simulation data and its application to galaxy formation models. Definitions range from disruption below the resolution of the simulation [27], coalescence defined by multiple progenitor subhalos sharing a common descendant subhalo [12, 28], losing a given fraction of the initial specific orbital angular momentum [11, 13], or reaching a certain physical separation [35]. To further complicate matters, the choice of the (arbitrary) threshold used within these definitions can modify the measured merger rate. For instance, J. M. Solanes et al. [36] demonstrated that defining subhalo mergers based on a threshold of 5% of the initial specific angular momentum leads to the premature termination of tracked orbital decay. These inconsistencies are particularly critical for mergers driven by dynamical-friction sinking, because the choice of an arbitrary threshold can systematically bias the inferred physics. The variety of merger definitions in use reflects our insufficient understanding of a core question regarding the process: what defines a merger physically? Addressing this question can provide a more self-consistent and physically motivated definition of merger events.
Another complication stems from the hierarchical nature of accretion itself. It is common for subhalos to be accreted as members of groups (“group infall”), which naturally establishes a nested, multi-level hierarchy within the host halo [37, 38, 39, 40, 41, 42]. This hierarchy entails that merging with the central subhalo is not the only fate for accreted subhalos, and some will merge with other satellite subhalos during their orbital evolution. Although central–satellite and satellite–satellite mergers are often analyzed separately, detailed studies that track the fate of accreted groups of subhalos and quantify the satellite–satellite merger rate remain limited.
By leveraging a dark-matter-only (DMO) cosmological simulation analyzed with the hbt+ subhalo finder [43, 44], we investigate subhalo mergers driven by dynamical friction. Unlike methods that rely on thresholds measured on the scale of the subhalo (e.g., bound particle number or orbital angular momentum), hbt+ determines the phase-space coalescence of a subhalo with the dense core of another subhalo. We refer to mergers identified through this physically motivated criterion as subhalo sinking (see Section 2.2 for details). The term “sinking” is used interchangeably with “merger” throughout this work, but specifically denotes those merger events identified by the hbt+ algorithm. Established terms such as “central–satellite merger” and “major merger” retain their conventional meanings. This work aims to:
  1. (i)
    Characterize the final fates of sunken subhalos within multi-level hierarchical systems.
  2. (ii)
    Introduce a systematic definition of the subhalo sinking rate that enables quantitative comparison across different hierarchical levels; and
  3. (iii)
    Examine how the sinking rate and radial location of sinking events depend on redshift, subhalo mass, and host halo mass.
Alhough the mergers among subhalos are rooted in the largely self-similar merger tree of the host halo [40], we will show that subhalo mergers are much more complex and deviates from self-similarity, due to additional physical processes during subhalo evolution since their accretion.
The paper is organized as follows. Section 2 describes the simulation as well as the hbt+ algorithm. Section 3 displays the fates of sunken subhalos within the multi-level hierarchy and classifies the cross‑level sinking pathways. Section 4 introduces two complementary definitions of the subhalo sinking rate—group‑level and halo‑level—and presents their dependence on mass ratio, redshift, and host halo mass. Section 5 examines the radial distribution of sinking events and decomposes it by sinking type. Section 6 discusses the resolution dependence of the core scales and the implications for defining merger criteria. It also examines the two mechanisms that drive subhalo sinking at the final stage of orbital decay. Section 7 summarizes our main findings.
2 Methods
2.1 Simulation & Subhalo Finder
Our analysis is based on a dark-matter-only simulation from the CosmicGrowth suite [45], specifically the L600 run. The simulation adopts a WMAP cosmology with parameters $H=71\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{M_{0}}=0.268$ and $\Omega_{\Lambda}=0.732$, and follows the evolution of $3072^{3}$ particles within a periodic box of side length $600\,\mathrm{Mpc}\,h^{-1}$. The particle mass is $m_{p}\approx 5.5\times 10^{8}\,\mathrm{M}_{\odot}\,h^{-1}$. The gravitational force is softened with a Plummer-equivalent softening length of $\epsilon=3.3\,\mathrm{kpc}\,h^{-1}$ in the comoving coordinate. In this work, halo masses are defined as the mass enclosed within the radius where its inner matter density equals the virial density predicted from the spherical collapse model, $\rho_{\rm vir}=\Delta_{\rm vir}\rho_{\rm crit}$, where $\Delta_{\rm vir}=18\pi^{2}+82[\Omega_{\rm m}(z)-1]-39[\Omega_{\rm m}(z)-1]^{2}$ [46].
We identify subhalos and construct their merger trees using the hbt+ (Hierarchical Bound-Tracing) algorithm [44]. In hbt+, the tracking algorithm operates in two distinct regimes: (1) when a subhalo is the central structure of its own host halo, its bound particles are continuously updated using the particle reservoir of the corresponding FoF group; (2) once it falls into a more-massive host and becomes a satellite, hbt+ records its particle membership at the infall time and subsequently removes particles that become unbound during its orbital evolution. A subhalo is considered resolved and is tracked as a self-bound object as long as it retains at least $N_{\mathrm{min}}=20$ bound particles. If the number of bound particles of a subhalo falls below this threshold, this subhalo is flagged as unresolved. hbt+ subsequently tracks it as an “orphan” subhalo, following only the position and velocity of the most bound particle when it was last resolved thereafter.
hbt+ provides a complete and self-consistent record of the hierarchical relationships between subhalos, tracing their origin to group accretion events. Within this structure, the hierarchical level (denoted by the symbol $\ell$ hereafter) of a subhalo is quantified by the number of distinct host subhalos that sequentially contain the subhalo. Specifically, the central subhalo is designated as $\ell=0$. Based on the hierarchical level, we define the following terminology to describe hierarchical relationships:
Child Subhalo: A subhalo (of level $\ell$) is considered the child of the specific subhalo of level $\ell-1$ or shallower levels, in which it is sequentially embedded.
Direct Parent Subhalo: The immediate host subhalo within which a child subhalo is directly embedded. This corresponds to the host subhalo at the next shallower hierarchical level.
Parent Subhalo: The set of all hosts that contain the child subhalo at successively shallower hierarchical levels. This includes its direct parent and all ancestors up to the central subhalo.
For example, a level-3 subhalo is embedded within a level-2 subhalo (its direct parent), which in turn is embedded within a level-1 subhalo. This entire nested structure ultimately resides within the central subhalo of the host halo. Thus, this example subhalo has three parents: the central subhalo, the level-1, and the level-2 subhalos.
2.2 Identifying Sunken Subhalos
The orbital decay of a subhalo within its host halo is driven by dynamical friction, which continuously extracts orbital energy and angular momentum, leading to a gradual inspiral and eventual coalescence [47]. To identify the moment at which two subhalos coalesce, hbt+ employs a phase-space distance criterion.
Figure 1: A resolved sinking event between two subhalos. The figure shows the evolution of the kinematic separation between a level-1 subhalo and its central subhalo. This level-1 subhalo was accreted into its host halo at redshift $z=0.36$ with a peak mass ratio of $m_{\rm peak}/M_{\rm vir}=0.35$. We also present the spatial and velocity dispersions of the parent subhalo ($\sigma_{x}$ and $\sigma_{v}$), along with the instantaneous spatial separation ($\Delta_{x}$) and relative velocity ($\Delta_{v}$) between the two subhalos. The blue shaded region marks the stage during which the level-1 subhalo has already sunk.
The core idea is to compare the phase-space separation between two bound subhalos to the local resolution scale near the center of the parent subhalo. This scale is estimated from the parent subhalo: we compute the spatial ($\sigma_{x}$) and velocity ($\sigma_{v}$) dispersion of its 20 most bound particles, which is intended to represent the subhalo core. The phase-space separation $\delta_{\rm s}$ between a child subhalo and its parent is then defined as:
$\delta_{\rm s}=\frac{|\mathbf{x_{\rm c}}-\mathbf{x}_{\rm p}|}{\sigma_{x}}+\frac{|\mathbf{v_{\rm c}}-\mathbf{v_{\rm p}}|}{\sigma_{v}},$ (1)
where $\mathbf{x_{\rm c}}$ and $\mathbf{v_{\rm c}}$ are the position and velocity of the most bound particle of the child, while $\mathbf{x_{\rm p}}$ and $\mathbf{v_{\rm p}}$ are the averaged position and velocity of the 20 most bound particles of the parent. A sinking event is flagged when $\delta_{\rm s}$ falls below a predefined threshold (set to 2 in this work), indicating the subhalo has become indistinguishable from a random particle in the core of the parent.
This check is performed iteratively at each snapshot for every subhalo, tracing its lineage upward from its direct parent toward the central subhalo of the host halo. We note that the algorithm only searches for sinking events between subhalos with a parent-child relationship (i.e., collectively accreted as a group of subhalos), not between peers at the same hierarchical level or unrelated substructures within the same host.
The original algorithm in hbt+ only considers a merger in the direction from a satellite subhalo to its parent. Following the approach of [48], we refine this sinking algorithm by evaluating the phase-space distance bidirectionally. Specifically, we compute:
$\delta_{\rm s,p}$ $=\frac{\lvert\mathbf{x}_{\rm c}-\mathbf{x}_{\rm p}\rvert}{\sigma_{x,\rm p}}+\frac{\lvert\mathbf{v}_{\rm c}-\mathbf{v}_{\rm p}\rvert}{\sigma_{v,\rm p}},$ (2)
$\delta_{\rm s,c}$ $=\frac{\lvert\mathbf{x}_{\rm p}-\mathbf{x}_{\rm c}\rvert}{\sigma_{x,\rm c}}+\frac{\lvert\mathbf{v}_{\rm p}-\mathbf{v}_{\rm c}\rvert}{\sigma_{v,\rm c}}.$ (3)
where $\delta_{\rm s,p}$ is identical to the original $\delta_{\rm s}$ in Equation 1, while $\delta_{\rm s,c}$ measures the distance from the parent to the child’s core. A sinking event is now identified when either of the two distances drops below the threshold of 2. When only $\delta_{\rm s,c}<2$ is satisfied, it typically corresponds to a child subhalo with a diffuse core, where its parent subhalo is embedded within the child’s own extended core. In all such cases, the child subhalo is still recorded as sinking into its parent.
Figure 1 shows the evolution of a representative level‑1 subhalo whose $\delta_{\rm s,p}$ value eventually falls below the detection threshold. Consistent with the case study in J. Han et al. [44], its evolution exhibits two distinct phases. After accretion, it experiences a rapid inspiral phase (2 to 3 orbital periods) driven by dynamical friction. This is followed by a stalling phase, where its spatial and velocity separations from the central subhalo cease decaying and instead oscillate around the central subhalo’s characteristic core scales, $\sigma_{x}$ and $\sigma_{v}$. This marks the onset of phase-space mixing with the central subhalo core at the given simulation resolution. Subsequent motion is dominated by random oscillations within the core rather than coherent inspiral. The oscillations could be caused by several factors, including particle resonances, the shape of the central density profile, and discreteness-induced fluctuations [49, 50, 51, 52]. The consistent two‑phase evolution indicates that the core‑mixing criterion implemented in the sinking algorithm provides a physically motivated diagnostic proxy for genuine coalescence in a dark‑matter‑only simulation at a given resolution. Reaching the $\delta_{\rm s,p}$ threshold typically corresponds to the end of the orbital‑decay phase.
Figure 2 further displays the evolution of a representative level‑1 sunken subhalo that is identified by the reverse‑link criterion. This subhalo also experiences a rapid inspiral phase lasting several orbital periods. Its $\delta_{\rm s,p}$, however, stalls at a value slightly above the threshold of 2. Meanwhile, near this point, the internal spatial dispersion of the subhalo’s own core grows substantially. This growth is driven by the tidal heating when approaching the center of the host halo and in turn continuously reduces $\delta_{\rm s,c}$. Generally, its evolution is still consistent with the two-phase nature and confirms that the refined sinking algorithm successfully identifies additional sinking events that the single‑distance ($\delta_{\rm s,p}$ only) criterion would miss.
Figure 2: Evolution of $\delta_{\rm s,p}$ and $\delta_{\rm s,c}$ for a subhalo that can be identified as sunken only via the $\delta_{\rm s,c}$ criterion (From the dedicated hbt+ run in Section 6.2). The red and blue solid lines trace the evolution of $\delta_{\rm s,c}$ and $\delta_{\rm s,p}$, respectively. The gray dashed line shows the evolution of the spatial dispersion ($\sigma_{x,\rm c}$) of the child subhalo’s core, while the gray solid line marks the threshold value of 2. The blue shaded region is used to represent the evolution after the subhalo sinking. The time axis begins at the snapshot when both the child subhalo and its direct parent are first recorded. The termination of the red line indicates that the child subhalo subsequently becomes unresolved.
A more detailed population‑level comparison between the subhalos identified by the refined algorithm and those found by the single‑side version is presented in Section 6.2, where the physical mechanisms of sinking are also discussed. Generally, this modification detects approximately 32% more sinking events than the original single-sided criterion in halos with $M_{\rm vir}>10^{12}\mathrm{M}_{\odot}\,h^{-1}$, regardless of the redshift. Due to tidal heating, the child subhalo usually quickly disrupts in a few dynamical times in the reverse-link case, so that omitting these sinking events does not have a significant influence on the final subhalo population. Nevertheless, we include these reverse-linked cases in our analysis for a more complete census of subhalo sinking.
Figure 3: Distributions of the levels of the final coalesced parent subhalo for child subhalos at a given level. The four panels correspond to child subhalos of the first four levels. Dark blue bars denote results in more massive halos, while the light blue bars denote those in less massive halos.
2.3 Subhalo Samples
To select a stable population of subhalos that are robust against rapid post-infall disruption, we apply a peak-mass cut
$m_{\rm peak}\geq 100m_{\rm p},$ (4)
where $m_{\rm p}$ denotes the particle mass. This criterion guarantees that the selected subhalos are well-resolved when they are accreted. Moreover, when a subhalo is stripped down to a single particle (becoming an unresolved subhalo), it no longer satisfies the core physical condition for dynamical friction—namely, that the sinking object must be significantly more massive than individual background particles. Consequently, the post-disruption trajectory of an orphan is numerically unphysical and is unlikely to sink. For this reason, we exclude orphan subhalos from our analysis. To achieve this, we impose a mass cut for all the identified sinking events: we consider only events where both the sunken child subhalo and its parent have a mass
$m_{\rm sink}\geq 20m_{\rm p},$ (5)
which is measured at the time when the sinking event happens.
After applying these criteria, we identify a total of 415,976 sinking events occurring in host halos of virial mass $10^{12},10^{13}]\mathrm{M}_{\odot}\,h^{-1}$ across all snapshots, and 68,997 events in hosts of mass $10^{13},10^{14}]\mathrm{M}_{\odot}\,h^{-1}$. This sample forms the basis for our subsequent statistical analysis.
3 Hierarchical Sinking Pathways of Subhalos
The prevailing framework in the literature classifies subhalo mergers into two broad channels: central–satellite and satellite–satellite mergers. While conceptually clear, this binary scheme does not fully capture the complexity inherent in hierarchically structured systems. In the subhalo mergers defined via sinking, both channels encompass a diversity of sinking pathways, as satellites themselves often host their own substructures, creating a nested hierarchy.
Figure 4: The relative abundance of the cross-level sink scenarios through four pathways. The plot shows the distribution for all of the deep-level ($\ell>1$) subhalo sinking events across all redshifts. We use bars with different colors to indicate the cross-level sink events measured from different host halo mass ranges.
To move beyond this simplified view and reveal the detailed fate of subhalos at each specific level of this hierarchy, we introduce a refined classification based on the precise hierarchical levels of the sunken subhalos. This approach allows us to disentangle the complex web of subhalo sinking and quantify how each sinking pathway depends on the level of a subhalo within the subhalo hierarchy.
Figure 3 shows the distribution of parent levels for sunken child subhalos binned by their own level $\ell_{\rm sub}$. Our analysis reveals a clear hierarchical sinking pattern. The vast majority of subhalo sinking events occur between a subhalo and its direct parent (level $\ell_{\rm sub}-1$). This is consistent with the findings in R. E. Angulo et al. [29], where they demonstrated that subhalos involved in satellite–satellite mergers are dynamically associated with each other prior to infall, rather than forming through random encounters within the dark matter halo.
However, cross-level sinking also constitutes a non-negligible fraction of sinking events, and this fraction increases slightly with $\ell_{\rm sub}$. For these cross-level events, the level difference between the child and its final parent is predominantly 2. Notably, we find that only a very small fraction ($\approx 0.5\%$) of sinking events occur directly between the central subhalos (level 0) and child subhalos with $\ell_{\rm sub}>2$. It indicates that nearly all the central-satellite mergers are contributed by the subhalos and sub-subhalos in the simulation. Furthermore, we find that this parent-level distribution exhibits little dependence on the mass of the main host halo, indicating that the underlying sinking process is governed by local hierarchy rather than global halo properties.
We classify these cross-level sinking events into four distinct pathways based on the evolutionary sequence of the child and its direct parent:
  • Overtaking Sinking: The child sinks into a shallower-level parent before its direct parent becomes a sunken subhalo. This could correspond to the case where a deeper-level subhalo becomes dynamically dissociated from its direct parent due to the tidal force from its host.
  • Trailing Sinking: The direct parent sinks first, and then the child sinks into a shallower-level parent.
  • Orphaned Sinking: The child sinks after its direct parent has been tidally disrupted. The occurrence of this pathway is sensitive to the simulation’s particle resolution. A higher resolution could potentially resolve the parent for longer, reducing the orphaned population.
  • Synchronous Sinking: Both the child and its direct parent sink into a shallower-level parent in the same snapshot. This pathway is sensitive to the time resolution of the simulation outputs, as a finer time sampling might separate these events across adjacent snapshots.
Subsequently, we measure the relative abundance of these four pathways in all the cross-level sinking events. The results are shown in Figure 4. Orphaned Sinking and Overtaking Sinking are the two dominant pathways in the cross-level sinking events. The high fraction of Orphaned Sinking suggests that insufficient particle resolution plays a significant role in facilitating cross-level events. The prevalence of Overtaking Sinking indicates that tidal stripping of the nearby massive subhalos can efficiently unbind a subhalo from its direct parent, allowing it to sink independently toward another one. It highlights the complex interplay of physical stripping and numerical effects in determining the fate of group-accreted subhalos [53, 54, 55, 56].
Our analysis of several pathways of the subhalo sinking has important applications in the semi-analytical model. In some models, when a subhalo group is accreted, the merger clocks of its satellites are reset based on the new host halo properties [57, 58]. Our findings offer a more nuanced view. The prevalence of next-level sinking supports the idea that many satellites remain bound to their group halo, for which maintaining the original merger timescale may be appropriate [59, 60].
4 The Subhalo Sinking Rate
The merger rates between subhalos and galaxies have been extensively studied in the literature [34, 31, 30, 27, 32, 28, 33]. In this section, we examine the subhalo merger rate identified via subhalo sinking and revisit subhalo sinking by introducing a finer classification based on the hierarchical level $\ell$ of satellites within the subhalo structure. We focus exclusively on sinking events between child subhalos and their direct parent subhalos, as these account for the vast majority of sinking events ($90\%$). We compare the sinking rates across different groups and halos and find that satellite–satellite sinking events constitute a non-negligible fraction of the overall sunken population.
4.1 Two Complementary Definitions of the Sinking Rate
Our finding that subhalo sinking occurs predominantly between a child and its direct parent motivates two complementary definitions:
1. Group-level sinking rate. A direct parent subhalo of level $\ell-1$ and its children of level $\ell$ can be defined as a subhalo group of $\ell-1$ within a dark matter halo. Investigating the subhalo sinking rate within each such group provides a quantitative measure of merger rate across hierarchical levels and reveals how the efficiency of internal orbital decay within a group depends on its hierarchical level. Here, we define the group-level sinking rate for subhalos of level $\ell$ as:
$f_{\mathrm{group},\ell}(\mu_{\rm rel}\mid m_{\mathrm{parent}},z)=\frac{\Delta N_{\mathrm{sink},\ell}(\mu_{\rm rel}\mid m_{\mathrm{parent}},z)}{N_{\rm group,\ell-1}\,\Delta\tau\,\Delta\ln\mu_{\rm rel}},$ (6)
where $\mu_{\rm rel}\equiv m_{\rm child}/m_{\rm parent}$ is the ratio of the child subhalo’s peak mass to the peak mass of the descendant of its direct parent after merging. We define $\Delta\tau\equiv\Delta t/t_{\rm dyn}(z)$, where $\Delta t$ is the time interval between two snapshots and $t_{\rm dyn}(z)$ is the dynamical timescale at $R_{\rm vir}$ at that redshift. The denominator is constructed as follows: at a given simulation snapshot (redshift $z$), we select all level-$(\ell-1)$ groups whose group mass—defined as the peak mass of the direct parent subhalo at that snapshot—falls within a specified range. $N_{\rm group,\ell-1}$ is the total number of such groups. The numerator $\Delta N_{{\rm sink},\ell}(\mu_{\rm rel}\mid m_{\mathrm{parent}},z)$ counts, within the same selected groups, all events occurring over the interval $\Delta\tau$ where a level-$\ell$ subhalo sinks into its direct level-$(\ell-1)$ parent, with a mass ratio residing within $\ln\mu_{\rm rel},\ln\mu_{\rm rel}+\Delta\ln\mu_{\rm rel}]$. This measures the sinking rate per subhalo group per unit time. While conceptually similar to A. R. Wetzel et al. [28], our key difference is the dissection of the rate by subhalo level $\ell$, allowing us to isolate distinct sinking channels (e.g., level-2 into level-1 vs. level-1 into the central subhalo) that are often blended in previous studies.
2. Halo-level sinking rate. This perspective aggregates all sinking events of level-$\ell$ subhalos within an entire host halo. We define the halo-level sinking rate for subhalos of level $\ell$ as:
$f_{\rm halo,\ell}(\mu_{\rm peak}\mid M_{\mathrm{vir}},z)=\frac{\Delta N_{{\rm sink},\ell}(\mu_{\rm peak}\mid M_{\rm vir},z)}{N_{\rm halo}\,\Delta\tau\,\Delta\ln\mu_{\rm peak}},$ (7)
where $\mu_{\rm peak}\equiv m_{\rm peak}/M_{\mathrm{vir}}$ is the ratio of the subhalo’s peak mass to the host halo’s virial mass. $N_{\rm halo}$ is the number of host halos in the specified virial mass range at redshift $z$. The numerator $N_{\mathrm{sink},\ell}$ counts, over the time interval $\Delta\tau$ between snapshots, all level-$\ell$ child subhalos that sink into their direct parents within those host halos with peak mass ratios residing within $\ln\mu_{\rm peak},\ln\mu_{\rm peak}+\Delta\ln\mu_{\rm peak}]$. This measures the total number of level-$\ell$ sunken subhalos per host halo per unit time, matching conventions such as R. E. Angulo et al. [29]. Finally, to capture sinking events occurring in the outskirts, we include all sinking events within subhalo groups defined via the host halo’s FoF membership.
Figure 5: Top: The sinking rate for subhalo groups at redshift 0. Lines with different colors and symbols indicate the sinking rate for subhalos of different levels. The shaded regions of different colors correspond to results from more massive groups, and their thickness indicates the uncertainty derived from the Poisson error on the mean. Here, the group mass is determined by the peak masses of the direct parents of these sunken subhalos. Middle: The same as the top panel but for redshift 1. Bottom: The same as the top panel but for redshift 1.9.
4.2 Key Results and Interpretation
Figure 5 presents the group-level sinking rate $f_{\rm group}$ measured from two group mass bins at three different redshifts. We find that the subhalo sinking often occurs between massive child subhalos ($\mu_{\rm rel}>0.1$) and their direct parents for $\ell=1,2$. This is expected because subhalos with large relative mass ratios typically experience stronger dynamical friction and sink more rapidly.
Figure 6: Top: The sinking rate per halo for child subhalos at different redshifts. We used colored symbols to represent the sinking rate of level-1 subhalos at a given peak mass ratio at three different redshifts. The shaded regions in each plot indicate the sinking rate measured from more massive halo samples ($M_{\rm vir}$) with error bars. The error bars correspond to the Poisson noise derived from the total count of sink subhalos in each peak mass ratio bin. Bottom: Similar to the top panel, but measured from level-2 subhalos.
Another key finding is that the amplitude of the level-2 sinking rate at $z=0$ is approximately a factor of four lower than that of level-1. However, the sinking rate of these two levels becomes comparable at high redshifts ($z=1$ and $z=1.9$). This suppression could be attributed to three possible factors:
  • Starvation of child subhalos. Once a satellite group first falls into a larger halo, it ceases to accrete new member subhalos. As a result, its member subhalo population may become largely exhausted by $z=0$, leading to a lower merger rate.
  • Tidal mass loss. As deeper level subhalos on average have a larger mass ratio relative to their host halo at infall time [40], they tend to experience stronger tidal mass loss rates [61]. In addition, deeper-level subhalos tend to be accreted earlier, thus experiencing tidal stripping for a longer time. As a result, deeper level subhalos are more likely to become unresolved at the final redshift of interest, given the finite mass resolution of the simulation. Since we only count resolved mergers in this work, the merger rate of level‑2 subhalos may thus be partly hidden by resolution limitations compared with that of level-1 subhalos.
  • Tidal heating. The tidal field of the host halo, as well as interactions with neighboring substructures, can heat the orbits of child subhalos within the subhalo group, delaying their orbital decay and effectively suppressing the merger rate.
Figure 6 shows the halo-level sinking rate $f_{\rm halo}$. The level-1 rate closely resembles $f_{\rm group}$, differing only in the mass definition of its group. However, the level-2 rate reveals a substantial population of low-mass-ratio sinking events (extending down to $\mu_{\rm peak}\sim 10^{-2}$). This behavior can be understood as a natural consequence of the underlying parent population. The halo-level sinking rate $f_{\rm halo,2}$ is essentially a convolution of two quantities: it combines the group-level merger rate, $f_{\rm group,2}$, with the peak mass function of level-1 subhalos, $g_{1}(\mu_{\rm peak})$. Since the peak mass function steeply favors low-mass objects, the convolution is weighted toward small parent masses. Thus, even though $f_{\rm group,2}$ itself favors high $\mu_{\rm rel}$ and contributes little below $\mu_{\rm rel}\sim 0.1$, the convolution with the abundant low-mass parent population drives $f_{\rm halo,2}$ to continue rising toward low $\mu_{\rm peak}$.
Figure 7: Normalized halo assembly history as a function of cosmic time. Colored lines illustrate the median mass assembly histories of halos in the mass range $10^{12},10^{13}]\mathrm{M}_{\odot}\,h^{-1}$ across various redshifts. $\hat{M}(\tau)$ denotes the halo mass normalized by its final mass $M(z_{0})$. Time is normalized to the dynamical timescale of the halos at their respective redshifts.
Furthermore, we find that the peak mass ratio distribution for level-1 sunken subhalos gradually shifts toward lower values with increasing redshift and host halo mass. This appears in contradiction with the self-similarity of the halo merger tree revealed by [62, 40], which predicts a universal progenitor subhalo peak mass function across redshift. This apparent discrepancy can be understood from the different orbital phases of the subhalos since their accretion. Because lower mass ratio subhalos are on average accreted earlier (see Figure 8 in [40]), they also tend to sink earlier. As a result, lower mass ratio sinking events are more likely to happen in recently-formed halos, i.e., those that have assembled a large fraction of their masses (and subhalos) over a relatively short timescale. As shown in Figure 7, high redshift halos indeed assembled their masses more recently compared to lower mass ones, in accordance with the two-phase halo formation theory [63]. Consequently, their sink rate distribution shifts systematically towards lower merger ratios. This trend is consistent with the results shown in Figure 6. The same explanation applies to the host mass dependence, as high mass halos are formed more recently.
Figure 8: Upper panels: The evolution of the cumulative halo-level sinking rate with respect to the redshift in host halos within $10^{12},10^{13}]\mathrm{M}_{\odot}\,h^{-1}$ above two specific peak mass ratios. Colored solid lines in each panel show the best fit to the functional form $\propto(1+z)^{\alpha}$, with $\alpha$ being the fitted parameter. We use the Poisson error on the cumulative counts to calculate the error bars. Lower panels: Similar to the upper panels, but for the sink rate measured from more massive halos.
Tracing the redshift evolution of $f_{\rm halo}$ above fixed mass-ratio thresholds (Figure 8), we find that $f_{\rm halo}$ increases steadily in the form of $\propto(1+z)^{\alpha}$ for the first three levels. The slope $\alpha$ varies across the hierarchical levels with $\alpha=0.13,0.61,0.82$ for $\ell=1,2,3$ separately. Notably, for level‑2 subhalos with $\mu_{\rm peak}>10^{-2}$, their sinking rate becomes comparable to that of central–satellite (level‑1) sinking events, highlighting the significant aggregate contribution of satellite–satellite mergers. It is expected that the $f_{\rm halo}$ with $\ell=3$ can also rival the central-satellite rate when less massive subhalos are resolved in the simulations with higher particle resolution. The steady rise of the halo-level sinking rate for deep-level subhalos toward lower mass thresholds reveals that satellite–satellite mergers become increasingly significant as we resolve subhalos with progressively smaller peak mass ratios.
4.3 Comparison to Previous Work
A. R. Wetzel et al. [28] report that satellite–satellite mergers constitute only a negligible fraction of the overall subhalo merger population compared with the central-satellite mergers. In contrast, we find a comparable satellite–satellite sinking rate in both Figures 5 and 8. This apparent discrepancy, however, stems from differences in definitions and sample selection, rather than indicating a fundamental inconsistency.
A. R. Wetzel et al. [28] quantify the contribution of satellite–satellite mergers by measuring their fraction among all merger events whose descendant subhalos lie within a fixed bound mass bin (e.g., $10^{12},10^{13}]\,\mathrm{M}_{\odot}\,h^{-1}$). They find that satellite–satellite mergers account for only $5\%$$15\%$ of the total. This result is a natural consequence of their sample selection: within the descendant mass bins they consider, central subhalos dominate the population, while satellites constitute only a small fraction. Consequently, most merger events involving these descendants are central–satellite mergers, and satellite–satellite mergers are inevitably subdominant. While this statistic usefully characterizes the relative abundance of satellite–satellite mergers in a descendant-selected sample, it does not directly measure how frequently subhalos of shallower levels receive merged subhalos from deeper levels, nor does it reveal how the merger rate varies across hierarchical levels.
Our group-level sinking rate adopts a similar descendant‑mass selection but takes a different normalization perspective. Rather than comparing counts to the total sunken population, we classify sinking events by the level $\ell$ of the sunken subhalo and normalize each class by the number of its parent groups, $N_{\rm group,\ell-1}$, i.e., the population that could host such an event. It directly quantifies the intrinsic sinking frequency for each hierarchical level within its own parent group and enables meaningful comparisons across different levels. Under this definition, the comparable group-level sinking rates of level-1 and level-2 subhalos at high redshift indicate that, for a given peak mass, a central subhalo is no more likely to undergo a merger than a level-1 satellite.
The halo-level sinking rate captures a fundamentally different statistic. It measures the total number of sinking events occurring within a host halo, aggregated across all hierarchical levels and normalized per halo. This definition is similar in spirit to the merger rate adopted by R. E. Angulo et al. [29]. In this perspective, the steady increase of the level-2 sinking rate toward lower $\mu_{\rm peak}$—in contrast to the decreasing trend of level-1—reflects the growing contribution from the abundant population of low-mass level-1 subhalo groups.
Together, these two complementary definitions demonstrate that satellite–satellite sinking events are not a negligible minority, but rather an important process that must be accounted for in any realistic model of subhalo evolution.
Finally, we emphasize that our analysis is intentionally limited to well-resolved subhalos to avoid numerical artifacts. Therefore, we do not provide a universal fitting formula here. A more complete and resolution-independent model must incorporate sinking events involving unresolved subhalos, for instance via robust merger-timescale prescriptions [15, 14, 59]. A unified model that accounts for all levels and resolution effects will be presented in future work.
5 The radial distribution of the subhalo sinking events
Figure 9: Stacked radial probability function of sinking events for host halos of mass $10^{12},10^{13}]\,\mathrm{M}_{\odot}\,h^{-1}$ at different redshifts. Radial distances are normalized by $R_{\rm vir}$ at merger. Profiles are constructed from subhalos satisfying Eq. 4 and 5, and normalized by the number density profile. Solid/dashed lines represent next-level/cross-level mergers. For level-3, dashed and dotted lines distinguish mergers with level-1 and central subhalos, respectively. Different panels correspond to different redshift bins.
Analyzing where subhalo sinking events occur provides crucial insights into where dynamical-friction-driven sinking is effective and where it is suppressed by multi-body interactions and tidal effects. Figure 9 presents the stacked radial profiles of sinking events occurring in the host halos within the mass bin $10^{12},10^{13}]\mathrm{M}_{\odot}\,h^{-1}$ across several redshift intervals. For each sinking event, the radial distance of the sunken subhalo is normalized by the virial radius of its host halo at the sinking moment, and the resulting stacked distribution is further normalized by the subhalo number density profile, which is measured from a specific snapshot using subhalos with peak mass greater than $100m_{\rm p}$. This normalization removes the baseline spatial distribution of the overall subhalo population, allowing us to isolate the preferred locations of sinking events. When classified by the nature of the sinking partner, the radial distributions of sinking events fall into two distinct categories.
Central–satellite sinking events, in which a subhalo sinks directly into the central (level-0) subhalo, exhibit a centrally concentrated radial profile regardless of the levels of subhalos. Whether the sinking involves a level-1, level-2, or level-3 subhalo sinking into the central, the distribution peaks at $r\leq 0.1R_{\rm vir}$ with an extended tail. The shape of this central concentration is remarkably consistent across levels, differing only in amplitude (see Figure 9). This universality reflects the common physical driver: dynamical friction operating deep within the host’s potential well. Notably, a small fraction of these central-satellite events also occur at large radii ($r\gtrsim 0.5R_{\rm vir}$). These distant events can be attributed either to central subhalos with unusually diffuse spatial distributions or, more commonly, to satellite subhalos whose assigned positions are associated with large core spatial dispersions ($\sigma_{R}$) via the reverse-link criterion. In both cases, the sunken subhalos have $\sigma_{R}$ values in the top 5% of the overall spatial dispersion distribution (Figure 11), confirming their anomalously diffuse nature.
[…]
The dominance of satellite–satellite sinking events in the outer regions is consistent with previous findings [29, 64]. This suggests that subhalo groups near the center cannot retain their member subhalos long enough for orbital decay to complete.
[…]
Finally, we find that the redshift evolution of the radial profiles is mild. The overall shapes and characteristic scales remain similar from $z=0$ to $z\approx 1.5$, suggesting a nearly universal radial preference for subhalo sinking across cosmic time.
Taken together, the radial probabilities highlight that the subhalo sinking is intimately connected to the hierarchical nature of structure formation. While central–satellite mergers dominate the innermost halo regions, satellite–satellite mergers preferentially occupy the outskirts, underscoring the key roles played by group infall, tidal perturbations, and the multi-level subhalo hierarchy in governing the locations of subhalo coalescence.
6 Discussion
6.1 Resolution Dependence
In this work, we employ the sinking algorithm to identify the physically resolved merger phase between subhalos within a given simulation resolution.
Figure 11: Upper panels: Spatial dispersion of the 20 most bound particles of subhalos at $z=0$ as a function of bound mass. The left and right panels correspond to central and satellite subhalos, respectively. The blue solid line indicates the median dispersion in each mass bin, and the black dashed line marks the simulation softening length. The region between the two gray dashed lines in each panel encloses the central 90% (5th–95th percentile) of subhalos at a given mass. For clarity, only $10^{5}$ randomly chosen subhalo samples are shown, with colors indicating subhalo density. Lower panels: Velocity dispersion of the 20 most bound particles of subhalos, following the same layout as the upper panels. The blue solid line represents the median of the overall velocity dispersion distribution, and the black dashed line shows the $\sigma_{v}-N_{\rm particle}$ fitting formula derived from the Millennium-II simulation in [44]. As above, the gray dashed lines enclose the 5th–95th percentile range in each mass bin.
This approach is motivated by the observed two‑phase evolution of subhalos: a rapid inspiral followed by a stalled, oscillatory regime near the parent core. To evaluate the robustness and applicability of this framework, we examine how the characteristic core scales ($\sigma_{x},\sigma_{v}$) depend on subhalo bound mass and vary across simulations with different resolutions.
Figure 11 shows the distribution of $\sigma_{x}$ and $\sigma_{v}$ as a function of subhalo bound mass for both central and satellite subhalos. The median spatial dispersion $\sigma_{x}$ decreases with subhalo mass and asymptotes to a value above the force softening length for subhalos containing at least 200 particles. This implies that, when applying this method, the minimum requirement is that the parent subhalo must be well-resolved (e.g., $M_{\rm vir}\geq 10^{12}\,\mathrm{M}_{\odot}\,h^{-1}$ or $m_{\rm parent}\geq 10^{12}\,\mathrm{M}_{\odot}\,h^{-1}$ in this work) for the core scales to be reliably measured. The velocity dispersion $\sigma_{v}$, in contrast, increases with mass, reflecting the depth of the gravitational potential. Notably, the median values of both distributions are nearly identical for satellites and centrals, indicating that these core scales are robust to tidal stripping until a subhalo becomes extremely disrupted. These results establish that the stalling scales of orbital decay are set by the subhalo’s intrinsic potential and the simulation’s numerical resolution.
Figure 12: Left: xyxy-plane projection of the orbital trajectory of a level-1 subhalo. The red solid line traces the trajectory of this level-1 subhalo (those directly orbiting the main halo) from its first crossing of the host’s virial radius to the present epoch. The main halo center at $z=0$ is marked by a black dot, and its virial radius ($R_{\rm vir,0}\approx 0.8$ Mpc $h^{-1}$, corresponding to a mass $M_{\rm vir,0}\approx 5.9\times 10^{13}\mathrm{M}_{\odot}\,h^{-1}$) is shown as a blue circle. The subhalo’s current position (at $z=0$) is indicated by a gold star, and its instantaneous velocity vector is denoted by an orange arrow. We use the orange-dashed circles to show the half-mass radius of this subhalo at the final spatial location during the evolution. Meanwhile, we also display the number of the bound particles at that time near the orange circles. At present, the subhalo resides at a normalized separation $\Delta r/R_{\rm vir}=0.3$ from the halo center, while its orbital angular momentum has decayed to only $\sim 3.4\%$ of the value it possessed when it initially crossed the virial boundary. Middle: yzyz-plane projection of the same subhalo. Right: Evolution of several orbital parameters of the level-1 subhalo, normalized by their values at the time of first infall. The black solid line shows the evolution of the specific angular momentum, while the red and blue lines denote the radial distance and tangential velocity, respectively, computed from the position and velocity averaged over bound particles. The intrinsic spin of the subhalo contributes little to the specific angular momentum during evolution.
The inherent dependence of $\sigma_{x}$ and $\sigma_{v}$ on numerical resolution becomes evident when comparing our results with the higher-resolution Millennium-II analysis in J. Han et al. [44], which adopts a softening length $\epsilon=1\,\mathrm{kpc}\,h^{-1}$ and a particle mass $m_{\rm p}=6.89\times 10^{6}\,\mathrm{M}_{\odot}\,h^{-1}$. In the $\sigma_{x}$ distribution, the median reported in J. Han et al. [44] asymptotes to approximately the softening length, whereas our median lies at about 1.5 times our softening length. This offset likely stems from differences in simulation codes, softening schemes, or the systematic shift induced by adopting different softening lengths. For $\sigma_{v}$, while the overall trends with subhalo mass are consistent with the fitting formula in J. Han et al. [44], our measured velocity dispersions are systematically larger. This is a direct consequence of particle mass resolution: for a given subhalo mass, the 20 most bound particles (from which $\sigma_{x}$ and $\sigma_{v}$ are computed) sample a larger region of phase space than in a higher-resolution simulation, due to the larger particle mass in our simulation. This naturally leads to a higher velocity dispersion according to the rising velocity dispersion profile in the inner region of dark matter halos. The resolution dependence underscores that a sinking event identified at a given resolution does not necessarily correspond to a genuine merger in the limit of infinite or even higher resolution. Nevertheless, within the constraints of the current simulation, these events serve as a well-motivated and physically diagnostic proxy for the coalescence of subhalo cores.
Moreover, the fact that stalling scales are set by subhalo-intrinsic properties, rather than by host halo scales, carries a key implication: it challenges merger criteria defined on scales comparable to or larger than the subhalo itself, particularly those relying on arbitrarily chosen static thresholds—such as a subhalo crossing $0.1R_{\rm vir}$ or losing a fixed fraction (e.g., 95%) of its initial specific angular momentum. The endpoint of orbital decay is not a universal fraction of the host virial radius $R_{\rm vir}$, but is instead set by the subhalo’s own core scale $\sigma_{x}$; the ratio $\sigma_{x}/R_{\rm vir}$ itself varies systematically across mass and resolution. Figure 12 illustrates this clearly: a subhalo in our simulation had lost 96% of its initial specific angular momentum yet remained at $0.3R_{\rm vir}$ from the host center. A fixed physical threshold, therefore, unavoidably over-merges systems of different masses and resolutions, which explains why orbital evolution often continues well beyond the merger times identified by such rules, as also noted by J. M. Solanes et al. [36].
6.2 Physical Mechanisms of Subhalo Sinking
6.2.1 Classification of Sinking Population
To assess whether a bidirectional sinking criterion recovers more physically coalesced subhalos than a single‑direction approach, we perform a dedicated hbt+ run on the L600 simulation with the MergeTrappedParticle option disabled. This configuration allows us to continue tracking $\delta_{\rm s,p}$ and $\delta_{\rm s,c}$ for sunken subhalos until they become unresolved, thereby enabling the identification of subhalos that satisfy only $\delta_{\rm s,c}<2$ during their evolution.
We first examine the distribution of satellite subhalos in the 2D plane of $\delta_{\rm s,p}$ and $\delta_{\rm s,c}$, following J. Han et al. [44]. Figure 13 shows this distribution for all resolved subhalos with respect to their direct parents in host halos of mass $M_{\rm vir,0}>10^{12}\,\mathrm{M}_{\odot}\,h^{-1}$ at $z=0$. The two marginalized distributions reveal a clear bimodality: one component consists of subhalos whose distance values cluster around tens, with the majority still orbiting and not yet phase‑mixed with their direct parents; the other component corresponds to sunken subhalos. This bimodal pattern persists at $z=1$, while at $z=1.9$ the sunken population is noticeably smaller.
Figure 13: The 2D density distribution of $\delta_{\rm s,p}$ and $\delta_{\rm s,c}$ of the resolved subhalos with respect to their direct parents at several specific redshifts. Here, we select satellites that are still resolved within the host halos with masses greater than $10^{12}\,\mathrm{M}_{\odot}\,h^{-1}$ at those redshifts. In each plot, the marginal density distributions of the $\delta_{\rm s,p}$ and $\delta_{\rm s,c}$ are also displayed on the upper and right panels separately. The horizontal and vertical dashed lines here represent the threshold value used to identify the sunken subhalos in this work.
To separate the subhalo population directly into sunken and orbiting components, we measure the minimum value of $(\delta_{\rm s,p},\delta_{\rm s,c})$ attained by each subhalo while it remains resolved. Our sample includes all subhalos hosted by FoF groups with parent halo masses $M_{\rm vir,0}>10^{12}\,\mathrm{M}_{\odot}\,h^{-1}$ at $z=0$, including those that become unresolved by $z=0$ but were resolved at earlier times. Figure 14 presents the distribution of these minimum distances for the $z=0$ population, shown separately for resolved, unresolved, and combined samples.
Figure 14: Distribution of the minimum phase-space distances $\min(\delta_{\rm s,p})$ and $\min(\delta_{\rm s,c})$ attained by sunken subhalos during their evolution while they are still resolved. Top left: Distribution for the subhalos that are still resolved at $z=0$. The dashed lines at $\min(\delta_{\rm s,p})=2$ and $\min(\delta_{\rm s,c})=2$ partition the population into three regions: Region 1 (both $\min\delta_{\rm s,p}$ and $\min\delta_{\rm s,c}<2$): mergers identifiable by both the original and the refined sinking algorithm. Region 2 ($\min\delta_{\rm s,c}<2\leq\min\delta_{\rm s,p}$): mergers identified only through the side of $\delta_{\rm s,c}$ (uniquely captured by the refined sinking algorithm). Region 3 ($\min\delta_{\rm s,p}<2\leq\min\delta_{\rm s,c}$): mergers identified only through the side of $\delta_{\rm s,p}$. Top right & Bottom panels: The same analysis for the unresolved and total subhalo population at $z=0$.
Figure 15: The evolution of $\delta_{\rm s,p}$ and $\delta_{\rm s,c}$ for three level-1 subhalos identified as sunken solely by the $\delta_{\rm s,p}$ criterion. The red and blue solid lines trace, respectively, the evolution of these two distances relative to the subhalo’s direct parent. Different line styles denote whether the related quantity is measured from the perspective of the sunken child or the parent. For reference, black and orange lines show the temporal evolution of the characteristic core scales $\sigma_{x}$ and $\sigma_{v}$ of the subhalos. The time axis is zeroed at the first snapshot where both the child and its direct parent are simultaneously resolved; a line terminates either when the corresponding (child or parent) subhalo later becomes unresolved, or when the simulation reaches $z=0$ (the present epoch).
The minimum‑distance distribution differs markedly from the distribution of current distances. For resolved subhalos, the distribution clearly separates into three distinct sub‑populations: a sunken component, and an orbiting component (defined by $\min\delta_{\rm s,p}\geq 2$ and $\min\delta_{\rm s,c}\geq 2$) that further splits into two groups—one with minimum distances clustered around a few tens, and another with minimum distances between approximately 2 and 10. This bimodality within the orbiting population reflects two distinct orbital stages: the subhalos from the former group are still on their first passage, while the latter have undergone at least one passage and been substantially altered by dynamical friction.
For unresolved subhalos, the minimum‑distance distribution shows only two sub‑populations. Its orbiting component spans values from 2 to 100 but peaks between 2 and 10, indicating that most unresolved subhalos become unresolvable at a stage when they are already strongly influenced by dynamical friction.
Based on the thresholds $\delta_{\rm s,c}=2$ and $\delta_{\rm s,p}=2$, we classify all sunken subhalos into three regions according to which criterion identifies them as sunken:
  • Region 1 (doubly‑identified sinking): For approximately 59.8% of the sinking events, both $\min\delta_{\rm s,p}$ and $\min\delta_{\rm s,c}$ lie well below the threshold of 2 (peak near $(0.71,0.73)$). These robust events are identified by both the original and refined algorithms, with only minor differences in the assigned sinking time.
  • Region 2 (child‑dispersion–driven sinking): Here, $\min\delta_{\rm s,c}<2\leq\min\delta_{\rm s,p}$. This population is uniquely recovered by the refined algorithm and accounts for about 24.2% of the total sunken population. It is mainly contributed by child subhalos that exhibit larger internal spatial dispersion ($\sigma_{x,c}$) relative to their parents (see Figure 2).
  • Region 3 (parent‑dispersion–driven sinking): A minority of events (16.0%) satisfy $\min\delta_{\rm s,p}<2$ while $\min\delta_{\rm s,c}\geq 2$, meaning they are identified solely by the $\delta_{\rm s,p}$ criterion. This typically occurs when the parent subhalo has a slightly larger internal velocity or spatial dispersion than its child, causing $\delta_{\rm s,p}$ to become the smaller distance and fall just below the threshold. Representative evolutionary tracks for this sinking type are shown in Figure 15.
Overall, the refined algorithm recovers approximately 32% more sinking events than the original single‑direction version, with the addition dominated by child‑dispersion–driven sinking.
6.2.2 Driving Mechanisms: Orbital Decay versus Tidal Heating
We have shown representative case studies for Regions 1, 2, and 3 in Figures 12, and 15, respectively. Specifically, in reverse‑link cases, tidal heating appears to be the dominant mechanism driving the sinking process, causing $\delta_{\rm s,c}$ to drop below the threshold. In contrast, parent‑dispersion–driven and doubly identified sunken subhalos show less evidence of significant heating; their cores remain compact throughout the sinking process.
To statistically quantify the prevalence of tidal heating, we measure the ratio of spatial dispersion before and after each sinking event, defined as $\langle\sigma^{2}_{R,{\rm pre}}\rangle/\langle\sigma^{2}_{R,{\rm post}}\rangle$. For each event, we compute the average squared spatial dispersion over the three snapshots prior to sinking and over the three snapshots spanning the sinking moment and its immediate aftermath. The results are shown in Figure 16.
In child‑dispersion–driven cases, the majority ($70\%$) of sunken satellite subhalos experience significant tidal heating, with their $\sigma^{2}_{R}$ values increasing by a factor of $\sim 2$ after sinking, while their parent subhalos remain nearly intact and compact. This confirms that reverse‑link sinking events are primarily driven by heating at the interface between the two evolutionary phases. A direct consequence of the substantial tidal heating is that the child subhalo quickly disrupts in about 1-2 dynamical times, which happens to the majority of the child-dispersion-driven cases.
Conversely, in parent‑dispersion–driven and doubly identified sinking events, the sunken satellites remain compact throughout. Although the spatial dispersion of their parent subhalos may expand slightly, the child subhalos themselves show no evidence of tidal heating. In these cases, the transition between the two‑stage evolution is driven primarily by orbital decay rather than by heating.
These findings have important implications for the development of subhalo finders aimed at identifying subhalo mergers. For example, the newly developed subhalo finder Bloodhound [65] traces subhalo disruption through variations in the structural parameter $c_{\rm V}$ and claims that subhalo mergers are captured as part of these disruption events. Our results suggest that relying exclusively on this parameter may be insufficient to identify all potential subhalo mergers, particularly those driven by parent‑dispersion or identified through the doubly linked criterion, which do not exhibit the same heating signatures and are more persistent in subsequent mass evolution.
7 Summary and Conclusions
In this study, we have focused on the common merger pathway associated with subhalo sinking—orbital decay driven by dynamical friction. We use the term “sinking” synonymously with “merger”, both referring exclusively to the dynamical‑friction‑driven merger. The hbt+ subhalo finder identifies such sinking events via the phase-space coalescence of two well-resolved subhalo cores, a criterion motivated by the two-phase evolution of subhalos: a rapid inspiral followed by a stalled regime near the parent core, where orbital decay does not proceed indefinitely. The kinematic separation between satellites and their parents asymptotically approaches the parent’s core scales, $\sigma_{x}$ and $\sigma_{v}$, measured from its 20 most bound particles. As shown in Figures 1 and 2, orbital decay stalls once satellites reach these scales, producing resolution-dependent “core stalling” consistent with theoretical expectations. This framework enables us to trace dynamical-friction-driven mergers across multiple hierarchical levels. Our main findings are summarized as follows:
  1. 1.
    Most subhalos sink into their direct parents, but cross-level pathways exist. In Figure 3, we find that deep-level subhalos predominantly sink into their direct parents. However, tidal stripping, group infall, and resolution-driven disruption naturally generate several cross-level sinking pathways.
  2. 2.
    Resolved mergers are predominantly major mergers. Figure 5 reveals that sunken subhalos identified by the group-level rate $f_{\rm group}$ are predominantly major mergers, with $\mu_{\rm rel}>0.1$, reflecting the strong dynamical friction acting on massive infalling satellites. This preference for major mergers persists across hierarchical levels and redshifts.
  3. 3.
    The merger ratio distribution depends on the dynamical age of the host halo. The average merger ratio is lower in dynamically younger systems, corresponding to halos at a higher redshift or with a larger mass (Figure 6). This is because subhalos with a lower mass ratio are expected to merge earlier due to their earlier accretion time. Consequently, dynamically older halos are more likely to have exhausted their lower mass ratio mergers.
  4. 4.
    Sinking statistics depend strongly on the hierarchical level. Although deep-level subhalos typically have small peak-mass ratios relative to their host halos, they may still be massive relative to their direct parents. As a result, satellite-satellite mergers can contribute comparably to, or even exceed, the central-satellite merger rate at the low mass end (Figures 6 and 8).
  5. 5.
    Satellite–satellite mergers preferentially occur in the halo outskirts. In contrast to central–satellite mergers, which occur near the halo center, Figure 9 shows that sinking events of the deeper-level subhalos preferentially occur at larger radii.
  6. 6.
    Orbital decay and tidal heating are the two main mechanisms at the final stage of subhalo sinking. Child subhalos trapped in the core of their parent comprise the majority of the sunken population, primarily driven by dynamical friction. An additional $\sim 30\%$ of sunken subhalos can be identified as the parent subhalo trapped in the core of its child, primarily driven by tidal heating of the child subhalo.
Figure 16: Distribution of the ratio of spatial dispersions measured before and after sinking. The left panel shows the variation for child subhalos during the sinking process, while the right panel shows the corresponding variation for parent subhalos. Colored solid lines distinguish sinking events occurring in the three regions defined in the text. The sample includes all sinking events from halos with $M_{\rm vir}\geq 10^{12}\mathrm{M}_{\odot}\,h^{-1}$, regardless of redshift.
These results establish a coherent picture in which subhalo sinking in cosmological simulations is governed jointly by dynamical friction, hierarchical accretion, and numerical resolution, providing valuable insights for interpreting the evolution of satellite systems and for informing semi-analytic galaxy formation models with physically motivated multi-level sinking pathways. Unlike the self-similar host halo merger tree, these subhalo mergers break away from self-similarity due to the additional physical processes involved.
A caveat of the current analysis is that the sinking identification is conditioned on a given numerical resolution, due to the dependence of the subhalo core size on the resolution. In addition, subhalos that go below the mass resolution are not studied either. One way to overcome these limitations is to calibrate a physical equation for dynamical friction, likely coupled to the process of subhalo mass loss [61, 66]. In the current work, we have focused on the statistics of the mergers relative to the final (peak) mass of the host halo. It would be more advantageous to work with the parameters of the merger pair defined when the subhalo is first accreted into its direct parent, which are more relevant to the orbital evolution of the subhalo. We leave these investigations to future work. Nevertheless, the current analysis paves the way towards physical understanding and unified modelling of the hierarchical mergers among subhalos and satellites.
[…]
This work is supported by National Key R&D Program of China (2023YFA1607800, 2023YFA1607801), NSFC(12595312,12595311,12133006), 111 project (No. B20019), and the science research grants from the China Manned Space Project (No. CMS-CSST-2025-A04). We acknowledge the sponsorship from the Yangyang Development Fund. This project is supported in part by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 24DX1400100, ZJ2023-ZD-001). ZZ.L. acknowledges the Fundamental Research Funds for the Central Universities (KG202502). VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025). The computation of this work is done on the Gravity supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
[…]
The authors declare that they have no conflict of interest.